# Cooper-pair splitters as circuit elements for realizing topological superconductors

### v1

@@ Acknowledgements.
[…]
@@
Figure 7: Position labels $x$ for sites along the four wire legs $1,2,3$, and $4$ (green, orange, blue, and red, respectively). Each coordinate lies in the range $x=0,\ldots,L+1$ and the four wires cross at $x=L+1$.
@@
We find $L-1$ solutions for $q_{m}$ with $m=1,\ldots,L-1$ where $q_{m}$ is real. Additionally, there are two imaginary solutions
@@
with energies ±$\pm E_{\text{max}}=\pm 2t\cosh(\kappa)$ and $\kappa$ real. (These energies correspond to the lowest and largest eigenvalues shown in Fig. 8.) The wavefunctions for these states decay exponentially in magnitude as one moves away from the crossing. For $L$ even there is a zero energy solution, labeled by $m_{0}$, such that $q_{m_{0}}=\pi/2$ satisfies both Eq. (17) and (19). To summarize, we have accounted for all $14L+$ 1 modes by solving explicitly for $3L$ nodal and $L+1$ anti-nodal solutions.
Figure 8: Spectrum of the Hamiltonian in Eq. (7) for $\Delta=0$ and $\mu=0$. Pattern of nodal and anti-nodal flat bands, depicted in red and blue respectively for $\gamma=0$ when (a) $L$ is even and (b) $L$ is odd. The states at extreme ends of the spectrum have energy ±$\pm E_{\text{max}}$, are depicted in green, and correspond to anti-nodal states with imaginary momentum Eq. (20). The degeneracy for each band (per spin) is indicated with an x. The figure also show the splitting when inter X-molecule hopping $\gamma\neq 0$, with Chern numbers of each band indicated on the right for $\phi=\pi/2$. The bands with imaginary momentum (green) have an exponentially small splitting (in $L$) in the presence of $\gamma$, and the individual bands also present non-trivial Chern numbers $1,0,-1$, similar to other anti-nodal clusters (the splitting in the figure is not shown for clarity).
Figure 9: Wave-function configurations for (a) anti-nodal and (b) nodal orbitals. $\pm$ denotes the relative sign, up to a gauge choice, of the wave function on each leg. The colors green, orange, blue, and red label, respectively, the $1,2,3,$ and $4$ legs of the X-molecules. (c) The matrix elements $(\alpha_{\triangleright})_{rs}$ and $(\alpha_{\triangleleft})_{rs}$ are constructed from the hopping between $\psi^{(r,n)}$ and $\psi^{(s,n)}$ orbitals along the three sub-lattice sites $1,2$, and $3$ in the kagome lattice. The crossings are labeled by the same numbers as the sub-lattice sites 1, 2, and 3.
@@
As an example of how to construct the $\alpha_{\triangleright}$ and $\alpha_{\triangleleft}$ matrices, consider the hopping between the zero energy orbitals $\psi^{(1,n_{0})}$ and $\psi^{(2,n_{0})}$ orbitals, as defined in Eq. (13) with $n_{0}=(L+1)/2$, which accounts for the matrix elements $(\alpha_{\triangleright})_{12}$ and $(\alpha_{\triangleleft})_{12}$. As shown in Fig. 9(c), $(\alpha_{\triangleright})_{12}$ is the result of hopping between leg 3 (blue) of molecule 1 and leg 4 (red) of molecule 2. From Fig. 9(b) we see that orbital $\psi^{(1,n_{0})}_{3}$ has amplitude $-\mathcal{A}_{L}$ and $\psi^{(2,n_{0})}_{4}$ has amplitude $-\mathcal{A}_{L}$, so their product is
@@
Upon diagonalization of $H_{\mathbf{k}}^{\operatorname{eff},\operatorname{N}}$ we observe that a small $\gamma$ splits the 9 bands into two sets of three-fold degenerate flat bands with zero Chern number at energies ±$\pm 4\Gamma$, where $\Gamma=\gamma\mathcal{A}_{L}^{2}$, and three other bands sandwiched between them and separated by a gap (Fig. 10). We also observe that the eigenvectors in the middle bands consist only of superpositions of the orbitals $\psi^{(2,n_{0})}_{j}$ in Eq. (13), where opposite legs have the same sign. Keeping only these orbitals in the Hamiltonian $H_{\mathbf{k}}^{\operatorname{eff},\operatorname{N}}$ gives us:
@@
Figure 10: Band structure of $H_{\mathbf{k}}^{\operatorname{eff},\operatorname{N}}$, with $\Gamma=\gamma\,\mathcal{A}_{L}^{2}$. The flat bands at energies ±$\pm 4\Gamma$ are 3-fold degenerate. The three bands in the middle correspond to the effective model in Eq. (1).
The three bands in the middle carry Chern numbers $+1$, 00 and $-1$. Thus we conclude that these three bands reproduce the effective Hamiltonian per spin $h_{\mathbf{k}}$ in the effective theory Eq. (1). Note that since the 3-fold degenerate flat bands at ±$\pm 4\Gamma$ have vanishing Chern numbers, they do not contribute to the Hall conductivity at half-filling.
Our previous derivation indicates that regardless of the parity of $L$, ensured one is at half filling $\mu=0$, the low-energy physics is always determined by the effective Hamiltonian in Eq. (1). Not surprisingly, we find non-trivial topological numbers for appropriate values of flux. Since the physics in the middle bands are determined by the effective model, Chern numbers assume non-vanishing values for $\phi$ around half-integer multiples of $\pi$, $\phi_{p}=(p-1/2)\pi$ for $p\in\mathbb{Z}$, with opposite sign depending whether $p$ is either odd or even. The pattern of Chern numbers for all the bands is displayed in Fig. 8 (a) and (b) for $\phi=\pi/2$, where we note that the degenerate flat bands in the nodal clusters have vanishing Chern-number. As one moves away from such special flux values $\phi_{p}$, by an amount ±$\sim\pm\pi/4$, the bands acquire enough curvature to start crossing with each other. The electrons at the Fermi sea then acquire dispersion and present a metallic behavior, signaling a phase transition to the non-topological regime. Finally upon turning on superconductivity $\Delta\neq 0$, if we ensure that the chemical potential obeys $-|\Delta|<\mu<|\Delta|$, then a topological superconducting phase is achieved.
@@
The hopping amplitude for the effective model with one site per X-molecule can be seen from the above discussion to be $\Gamma=\gamma\mathcal{B}_{m_{0},L}^{2}$ or $\gamma\mathcal{A}_{L}^{2}$ depending on whether $L$ is even or odd, respectively. Both the amplitudes $\mathcal{B}_{m_{0},L}$ and $\mathcal{A}_{L}$ scale with the inverse of the square root of ${L}$. Thus, for both $L$ even and odd $\Gamma\sim\gamma/L$. In the regime in which the tunnel barriers between the X-molecules are small, i.e., when they are almost transparent so as to favor the case shown in Fig. 1(a), we can use $\gamma\sim t$, so that $\Gamma\sim t/L\sim v_{F}/\ell$. Finally, using the definition of the superconducting coherence length $\xi=v_{F}/\Delta$, we can express
@@
Supplementary Material B Josephson junction limit

### v2

@@ Appendix B Tight-binding model for wire network
$H_{t}$ $=-t\sum_{\mathbf{r}}\sum_{a=1,2,3}\;\sum_{\sigma=\uparrow,\downarrow}\;\sum_{\langle i,j\rangle\in\text{X}}\left(\;c^{\dagger}_{\mathbf{r},a,i,\sigma}\,c^{\,}_{\mathbf{r},a,j,\sigma}+\text{h.c.}\right)\;,$ (11)
@@
$H_{\mu}$ $=-\mu\sum_{\mathbf{r}}\sum_{a=1,2,3}\;\sum_{\sigma=\uparrow,\downarrow}\;\sum_{i\in\text{X}}\left(c^{\dagger}_{\mathbf{r},a,i,\sigma}\,c_{\mathbf{r},a,i,\sigma}+\text{h.c.}\right)\;,$ (13)
$H_{\Delta}$ $=\sum_{\mathbf{r}}\sum_{a=1,2,3}\;\sum_{i\in\text{X}}\left(\Delta_{a,i}\;c^{\dagger}_{\mathbf{r},a,i,\uparrow}\,c^{\dagger}_{\mathbf{r},a,i,\downarrow}+\text{h.c.}\right)\;,$ (14)
@@
Figure 13: Position labels $x$ for sites along the four wire legs $1,2,3$, and 44 (green, orange, blue, and red, respectively). Each coordinate lies in the range $x=0,\ldots,L+1$ and the four wires cross at $x=L+1$.
@@
We find L1L-1 solutions for $q_{m}$ with $m=1,\ldots,L-1$ where $q_{m}$ is real. Additionally, there are two imaginary solutions
@@
with energies $\pm E_{\text{max}}=\pm 2t\cosh(\kappa)$ and $\kappa$ real. (These energies correspond to the lowest and largest eigenvalues shown in Fig. 14.) The wavefunctions for these states decay exponentially in magnitude as one moves away from the crossing. For $L$ even there is a zero energy solution, labeled by $m_{0}$, such that $q_{m_{0}}=\pi/2$ satisfies both Eq. (20) and (22). To summarize, we have accounted for all $4L+1$ modes by solving explicitly for $3L$ nodal and $L+1$ anti-nodal solutions.
Figure 14: Spectrum of the Hamiltonian in Eq. (10) for $\Delta=0$ and $\mu=0$. Pattern of nodal and anti-nodal flat bands, depicted in red and blue respectively for $\gamma=0$ when (a) $L$ is even and (b) $L$ is odd. The states at extreme ends of the spectrum have energy $\pm E_{\text{max}}$, are depicted in green, and correspond to anti-nodal states with imaginary momentum Eq. (23). The degeneracy for each band (per spin) is indicated with an x. The figure also show the splitting when inter X-molecule hopping $\gamma\neq 0$, with Chern numbers of each band indicated on the right for $\phi=\pi/2$. The bands with imaginary momentum (green) have an exponentially small splitting (in $L$) in the presence of $\gamma$, and the individual bands also present non-trivial Chern numbers 1,0,11,0,-1, similar to other anti-nodal clusters (the splitting in the figure is not shown for clarity).
Figure 15: Wave-function configurations for (a) anti-nodal and (b) nodal orbitals. $\pm$ denotes the relative sign, up to a gauge choice, of the wave function on each leg. The colors green, orange, blue, and red label, respectively, the $1,2,3,$ and 44 legs of the X-molecules. (c) The matrix elements $(\alpha_{\triangleright})_{rs}$ and $(\alpha_{\triangleleft})_{rs}$ are constructed from the hopping between $\psi^{(r,n)}$ and $\psi^{(s,n)}$ orbitals along the three sub-lattice sites $1,2$, and $3$ in the kagome lattice. The crossings are labeled by the same numbers as the sub-lattice sites 1, 2, and 3.
@@
As an example of how to construct the $\alpha_{\triangleright}$ and $\alpha_{\triangleleft}$ matrices, consider the hopping between the zero energy orbitals $\psi^{(1,n_{0})}$ and $\psi^{(2,n_{0})}$ orbitals, as defined in Eq. (16) with $n_{0}=(L+1)/2$, which accounts for the matrix elements $(\alpha_{\triangleright})_{12}$ and $(\alpha_{\triangleleft})_{12}$. As shown in Fig. 15(c), $(\alpha_{\triangleright})_{12}$ is the result of hopping between leg 3 (blue) of molecule 1 and leg 4 (red) of molecule 2. From Fig. 15(b) we see that orbital $\psi^{(1,n_{0})}_{3}$ has amplitude $\mathcal{A}_{L}$ and $\psi^{(2,n_{0})}_{4}$ has amplitude $\mathcal{A}_{L}$, so their product is
@@
Upon diagonalization of $H_{\mathbf{k}}^{\operatorname{eff},\operatorname{N}}$ we observe that a small $\gamma$ splits the 9 bands into two sets of three-fold degenerate flat bands with zero Chern number at energies $\pm 4\Gamma$, where $\Gamma=\gamma\mathcal{A}_{L}^{2}$, and three other bands sandwiched between them and separated by a gap (Fig. 16). We also observe that the eigenvectors in the middle bands consist only of superpositions of the orbitals $\psi^{(2,n_{0})}_{j}$ in Eq. (16), where opposite legs have the same sign. Keeping only these orbitals in the Hamiltonian $H_{\mathbf{k}}^{\operatorname{eff},\operatorname{N}}$ gives us:
@@
Figure 16: Band structure of $H_{\mathbf{k}}^{\operatorname{eff},\operatorname{N}}$, with $\Gamma=\gamma\,\mathcal{A}_{L}^{2}$. The flat bands at energies $\pm 4\Gamma$ are 3-fold degenerate. The three bands in the middle correspond to the effective model in Eq. (1).
The three bands in the middle carry Chern numbers +1+1, 0 and 1-1. Thus we conclude that these three bands reproduce the effective Hamiltonian per spin $h_{\mathbf{k}}$ in the effective theory Eq. (1). Note that since the 3-fold degenerate flat bands at $\pm 4\Gamma$ have vanishing Chern numbers, they do not contribute to the Hall conductivity at half-filling.
Our previous derivation indicates that regardless of the parity of $L$, ensured one is at half filling $\mu=0$, the low-energy physics is always determined by the effective Hamiltonian in Eq. (1). Not surprisingly, we find non-trivial topological numbers for appropriate values of flux. Since the physics in the middle bands are determined by the effective model, Chern numbers assume non-vanishing values for $\phi$ around half-integer multiples of $\pi$, $\phi_{p}=(p-1/2)\pi$ for $p\in\mathbb{Z}$, with opposite sign depending whether $p$ is either odd or even. The pattern of Chern numbers for all the bands is displayed in Fig. 14 (a) and (b) for $\phi=\pi/2$, where we note that the degenerate flat bands in the nodal clusters have vanishing Chern-number. As one moves away from such special flux values $\phi_{p}$, by an amount $\sim\pm\pi/4$, the bands acquire enough curvature to start crossing with each other. The electrons at the Fermi sea then acquire dispersion and present a metallic behavior, signaling a phase transition to the non-topological regime. Finally upon turning on superconductivity $\Delta\neq 0$, if we ensure that the chemical potential obeys $|\Delta|<\mu<|\Delta|$, then a topological superconducting phase is achieved.
@@
The hopping amplitude for the effective model with one site per X-molecule can be seen from the above discussion to be $\Gamma=\gamma\mathcal{B}_{m_{0},L}^{2}$ or $\gamma\mathcal{A}_{L}^{2}$ depending on whether $L$ is even or odd, respectively. Both the amplitudes $\mathcal{B}_{m_{0},L}$ and $\mathcal{A}_{L}$ scale with the inverse of the square root of L{L}. Thus, for both $L$ even and odd $\Gamma\sim\gamma/L$. In the regime in which the tunnel barriers between the X-molecules are small, i.e., when they are almost transparent so as to favor the case shown in Fig. 1(a), we can use $\gamma\sim t$, so that $\Gamma\sim t/L\sim v_{F}/\ell$. Finally, using the definition of the superconducting coherence length $\xi=v_{F}/\Delta$, we can express
@@
Appendix C Josephson junction limit